# SAM Decoding: Speculative Decoding via Suffix Automaton

### v2

@@ 5 Related Work
Speculative Decoding. Speculative decoding is an approach that can significantly speed up large language models (LLMs) without compromising the quality of their outputs. This is achieved by increasing the number of tokens produced in each inference step while simultaneously decreasing the total number of inference steps required. The majority of speculative decoding techniques rely on smaller neural networks to create initial drafts during the inference process. These techniques are referred to as model-based speculative decoding methods. Early implementations of model-based speculative decoding, such as those described in Speculative Decoding Leviathan et al. (2023), primarily focused on generating draft sequences using pre-existing, smaller-scale LLMs. Subsequently, advancements like Medusa Cai et al. (2024) and SpecInfer Miao et al. (2024) introduced tree-based speculative methods and initiated the development of dedicated draft models tailored for speculative decoding. For instance, Medusa developed a set of specialized decoding heads designed to serve as draft models. More recently, EAGLE Li et al. (2024b, a) highlighted the significance of incorporating the hidden state of the model’s output into the speculative decoding process. In line with this insight, EAGLE trained a draft model that integrates the hidden state information from the model’s output, leading to more efficient acceleration across a range of tasks.
In contrast to model-based methods, certain approaches focus on generating drafts by leveraging the model’s recent prediction outcomes, such as the output from the decoding head. Examples of such methods include Lookahead Decoding Fu et al. (2024) and Token Recycling Luo et al. (2024). Lookahead Decoding keeps track of the $n$-grams of the generated text, whereas Token Recycling maintains the top $k$ probable next tokens for each tokens.
Moreover, some strategies generate drafts directly through the retrieval of previously generated texts or text corpora, utilizing $n$-gram matching. Notable among these are PLD Saxena (2023) and REST He et al. (2024). Our proposed SAM-Decoding method also employs retrieval for draft generation. However, it stands out due to its superior speed and accuracy, and it can be seamlessly integrated with other speculative decoding techniques.
@@
Efficient LLM Architecture. There is also work to improve the model’s inference speed from the perspective of model structure. This part of the work includes model distillation, quantization and pruning. Model distillation Sreenivas et al. (2024); Muralidharan et al. (2024) distills the knowledge of a large model into a small model, thereby speeding up inference while maintaining the model’s performance. Quantization Frantar et al. (2022); Xiao et al. (2023); Lin et al. (2024); Liu et al. (2024); Ashkboos et al. (2024b) reduces the number of bits required to store parameters and reduces the data transmission time from HBM to on-chip memory during inference, thereby achieving effective inference acceleration. Pruning Frantar and Alistarh (2023); Ashkboos et al. (2024a); Men et al. (2024); Chen et al. (2024); Hu et al. (2024); Sun et al. (2024); Zhang et al. (2024) is used to remove unimportant parameters in the model. For structured pruning, it can be combined with model distillation to train efficient small models, while semi-structured pruning can reduce the model’s memory access and computing overhead and improve the inference speed by combining special hardware, i.e., sparse tensor core.
@@ 6 Conclusion
In this work, we propose SAM-Decoding, an speculative decoding method via suffix automatons constructed from both generated text and text corpus. SAM-Decoding can efficiently retrieve drafts from retrieval sources, thereby accelerating inference. SAM-Decoding is also designed to seamlessly integrate with existing methods. Consequently, in scenarios where retrieval is not feasible, SAM-Decodingcan adaptively switch to alternative methods for draft generation. Experimental results demonstrate that when combined with state-of-the-art techniques, SAM-Decodingcan significantly enhance performance in multi-turn conversation, summarization, and retrieval-augmented generation tasks.
@@ A.1 Construction Process of Suffix Automaton
Algorithm 3 introduces the construction (Build-SAM) and expansion process (Expand) of Suffix Automaton, where INIT_SAM function will create a suffix automaton that only contains the root node. And for the root node, the link attribute value is $-1$, the next attribute value is empty, the length attribute value is 00, and the min_endpos attribute value is 00.
@@
A.2 Time Complexity of State Transfer
@@
Next, we analyze the upper bound of $\hat{c_{i}}$. Each state transition involves moving through the link edge zero or more times, followed by a move through the next edge. Transitioning through the link edge incurs a cost of 1 but decreases the potential by at least 1. Conversely, transitioning through the next edge also incurs a cost of 1 but increases the potential by 1. Consequently, the amortized cost $\hat{c_{i}}$ is bounded above by 2, leading to:
@@
Model Method Multi-turn Conversation Translation Summarization Question Answering Mathematical Reasoning Retrieval-aug. Generation #Mean Accepted Tokens Tokens/s Overall
Llama3-8B-Inst PLD 1.30$\times$ 1.12$\times$ 1.41$\times$ 1.03$\times$ 1.30$\times$ 1.53$\times$ $1.39$ 44.26 1.28$\times$
@@
SAM-Decoding[T] 2.09$\times$ 1.93$\times$ 2.04$\times$ 1.82$\times$ 2.32$\times$ 2.12$\times$ 2.63 85.73 2.05$\times$
EAGLE2 2.08$\times$ 1.95$\times$ 1.85$\times$ 1.80$\times$ 2.31$\times$ 1.87$\times$ 3.90 68.69 1.98$\times$
SAM-Decoding[E2] 2.23$\times$ 1.92$\times$ 1.93$\times$ 1.76$\times$ 2.23$\times$ 2.06$\times$ 3.91 70.40 2.03$\times$
@@
Model Method HumanEval MBPP HAGRID
#MAT Tokens/s Speedup #MAT Tokens/s Speedup #MAT Tokens/s Speedup
Llama3-8B-Inst PLD 1.30 42.39 1.18×1.18$\times$1.18 × 1.23 39.19 1.06×1.06$\times$1.06 × 1.50 45.15 1.34×1.34$\times$1.34 ×
Token Recycling 2.93 71.49 1.99×1.$9$ 2.89 74.05 2.02×2.02$\times$2.02 × 2.84 62.77 1.87×1.87$\times$1.87 ×
SAM-Decoding[T] 2.77 78.04 2.16×2.16$\times$2.16 × 2.65 76.76 2.09×2.09$\times$2.09 × 2.70 66.76 1.99×1.$9$
EAGLE2 4.74 85.58 2.37×2.37$\times$2.37 × 4.77 95.02 2.58×2.58$\times$2.58 × 3.97 63.30 1.86×1.86$\times$1.86 ×
SAM-Decoding[E2] 4.76 91.50 2.54×2.54$\times$2.54 × 4.34 89.15 2.43×2.43$\times$2.43 × 3.93 67.94 2.02×2.02$\times$2.02 ×
Table 5: Speedup of SAM-Decoding compared to the baselines on HumanEval, MBPP, and HAGRID.
Model Method Multi-turn Conversation Translation Summarization Question Answering Mathematical Reasoning Retrieval-aug. Generation #Mean Accepted Tokens Tokens/s Overall
@@
EAGLE2 3.10$\times$ 2.15$\times$ 2.58$\times$ 2.38$\times$ 3.19$\times$ 2.33$\times$ 4.42 56.06 2.63$\times$
@@
Model Method HumanEval MBPP HAGRID
#MAT Tokens/s Speedup #MAT Tokens/s Speedup #MAT Tokens/s Speedup
Vicuna-13B PLD 1.54 32.06 1.44×1.$4$ 1.40 29.43 1.30×1.30$\times$1.30 × 1.90 43.38 2.15×2.15$\times$2.15 ×
Token Recycling 2.79 46.03 2.07×2.07$\times$2.07 × 2.85 48.89 2.16×2.16$\times$2.16 × 2.90 40.97 2.03×2.03$\times$2.03 ×
SAM-Decoding[T] 2.79 50.87 2.28×2.28$\times$2.28 × 2.83 52.14 2.30×2.30$\times$2.30 × 2.99 48.33 2.40×2.40$\times$2.40 ×
EAGLE2 5.15 77.85 3.49×3.49$\times$3.49 × 5.26 82.13 3.63×3.63$\times$3.63 × 4.24 52.28 2.59×2.59$\times$2.59 ×
SAM-Decoding[E2] 4.97 73.78 3.31×3.31$\times$3.31 × 4.80 73.94 3.27×3.27$\times$3.27 × 4.41 56.17 2.78×2.78$\times$2.78 ×
Table 7: Speedup of SAM-Decoding compared to the baselines on HumanEval, MBPP, and HAGRID.
Model Method Multi-turn Conversation Translation Summarization Question Answering Mathematical Reasoning Retrieval-aug. Generation #Mean Accepted Tokens Tokens/s Overall
@@
Token Recycling 2.10$\times$ 1.84$\times$ 2.19$\times$ 1.88$\times$ 2.42$\times$ 1.92$\times$ 2.70 18.8 2.06$\times$
@@
EAGLE2 3.29$\times$ 2.31$\times$ 2.73$\times$ 2.51$\times$ 3.65$\times$ 2.46$\times$ 4.06 25.86 2.83$\times$
SAM-Decoding[E2] 3.36$\times$ 2.24$\times$ 2.88$\times$ 2.41$\times$ 3.43$\times$ 2.50$\times$ 4.08 25.60 2.81$\times$
@@
Model Method HumanEval MBPP HAGRID
#MAT Tokens/s Speedup #MAT Tokens/s Speedup #MAT Tokens/s Speedup
Vicuna-33B PLD 1.58 14.18 1.51×1.51$\times$1.51 × 1.34 11.82 1.30×1.30$\times$1.30 × - - -
Token Recycling 2.64 19.64 2.09×2.09$\times$2.09 × 2.76 20.92 2.19×2.19$\times$2.19 × - - -
SAM-Decoding[T] 2.73 22.44 2.39×2.39$\times$2.39 × 2.65 21.80 2.28×2.28$\times$2.28 × 2.60 19.74 2.26×2.26$\times$2.26 ×
EAGLE2 3.53 28.18 3.00×3.$0$ 5.04 37.35 3.91×3.91$\times$3.91 × 3.84 24.28 2.78×2.78$\times$2.78 ×
SAM-Decoding[E2] 3.61 29.49 3.14×3.14$\times$3.14 × 4.70 34.21 3.58×3.58$\times$3.58 × 3.79 24.45 2.80×2.80$\times$2.80 ×
Table 9: Speedup of SAM-Decoding compared to the baselines on HumanEval, MBPP, and HAGRID.
Model Method Multi-turn Conversation Translation Summarization Question Answering Mathematical Reasoning Retrieval-aug. Generation #Mean Accepted Tokens Tokens/s Overall
Vicuna-7B PLD 1.61$\times$ 1.56$\times$ 0.99$\times$ 2.47$\times$ 1.11$\times$ 1.51$\times$ 1.75 78.26 1.56$\times$
Token Recycling 2.08$\times$ 1.76$\times$ 1.97$\times$ 1.85$\times$ 2.35$\times$ 1.76$\times$ 2.82 98.39 1.96$\times$
@@
EAGLE2 2.66$\times$ 1.76$\times$ 2.18$\times$ 2.03$\times$ 2.63$\times$ 1.97$\times$ 4.34 110.56 2.21$\times$
@@
Table 10: Speedup of SAM-Decoding on A100 GPU compared to the baselines on Spec-Bench.
@@ Appendix B Additional Experiment Results
In this section, we present the results of the experiment on Llama3-8B-instruct.
Tables 4 and 5 present the speedup ratios of SAM-Decoding compared to baseline methods across the Spec-Bench, HumanEval, MBPP, and HAGRID datasets, utilizing the Llama3-8B-instruct model. In particular, SAM-Decoding , when paired with Token Recycling (SAM-Decoding[T]), brings speedups on all tasks and achieves the highest average speedup. Specifically, SAM-Decoding enhances the speedup ratio of Token Recycling from 1.92$\times$, 1.85$\times$, and 1.82$\times$ to 2.09$\times$, 2.04$\times$, and 2.12$\times$ for multi-turn conversation, summarization, and retrieval-augmented generation tasks, respectively. This improvement raises the overall speedup ratio of token recycling in the Spec-Bench dataset from 1.91$\times$ to 2.05$\times$. On the HumanEval and MBPP datasets, SAM-Decoding increases the speedup ratio of Token Recycling from 1.80$\times$ and 2.04$\times$ to 2.27$\times$ and 2.37$\times$, respectively. For the HAGRID dataset, the integration of SAM-Decoding with Token Recycling increases the speedup ratio from 1.87$\times$ to 1.99$\times$. Furthermore, beyond enhancing Token Recycling, SAM-Decoding also amplifies the performance gains of EAGLE2 in multi-turn conversation, summarization, retrieval-augmented generation, and context Q&A tasks. The speedup ratios were increased from 2.08$\times$, 1.85$\times$, 1.87$\times$, and 1.86$\times$ to 2.23$\times$, 1.93$\times$, 2.06$\times$, and 2.02$\times$ respectively.

### v3

@@ 5 Related Work
Speculative Decoding. Speculative decoding is an approach that can significantly speed up large language models (LLMs) without compromising the quality of their outputs. The majority of speculative decoding techniques rely on smaller neural networks to create drafts during the inference process. These techniques are referred to as model-based speculative decoding methods. Early implementations of model-based speculative decoding, such as those Speculative Decoding Leviathan et al. (2023), primarily focused on generating draft sequences using pre-existing, smaller-scale LLMs. Subsequently, advancements like Medusa Cai et al. (2024), SpecInfer Miao et al. (2024) and EAGLE Li et al. (2024b, a) introduced tree-based speculative methods and began the development of draft models tailored for speculative decoding.
In contrast to model-based methods, certain approaches focus on generating drafts through retrieval, utilizing $n$-gram matching, which we refer to the retrieval-based method. Notable among these are Lookahead Decoding Fu et al. (2024), PIAZhao et al. (2024), PLD Saxena (2023) and REST He et al. (2024). Token Recycling Luo et al. (2024), on the other hand, utilizes the previously generated token distribution to generate drafts, becoming a model-free method different from the retrieval-based method.
@@
Efficient LLM Architecture. There is also work to improve the model’s inference speed from the perspective of model structure. This part of the work includes model distillation, quantization and pruning. Model distillation Sreenivas et al. (2024); Muralidharan et al. (2024) distills the knowledge of a large model into a small model thereby speeding up inference while maintaining the model’s performance. Quantization Frantar et al. (2022); Xiao et al. (2023); Lin et al. (2024); Liu et al. (2024); Ashkboos et al. (2024b) reduces the number of bits required to store parameters and reduces the data transmission time from HBM to on-chip memory during inference. Pruning Frantar and Alistarh (2023); Ashkboos et al. (2024a); Men et al. (2024); Chen et al. (2024); Hu et al. (2024); Sun et al. (2024); Zhang et al. (2024) is used to remove unimportant parameters in the model. For structured pruning, it can be combined with model distillation to train efficient small models, while semi-structured pruning can reduce the model’s memory access and computing overhead and improve the inference speed by combining special hardware.
@@ 6 Conclusion
In this work, we propose SAM-Decoding, an speculative decoding method via suffix automatons constructed from both generated text and text corpus. SAM-Decoding can efficiently retrieve drafts from retrieval sources, thereby accelerating inference. SAM-Decoding is also designed to seamlessly integrate with existing SD methods. Consequently, in scenarios where retrieval is not feasible, SAM-Decoding can adaptively switch to alternative methods for draft generation. Experimental results demonstrate that SAM-Decoding outperform retrieval-based SD baselines. Meanwhile, when combined with state-of-the-art techniques, SAM-Decoding can significantly enhance their performance in Multi-turn Conversation, Summarization, Retrieval-augmented Generation, and Context Q&A tasks.
7 Limitation
On the one hand, as a retrieval-based speculative decoding method, the performance of SAM-Decoding depends on the task type as well as the quality of the retrieval source. Currently, we have collected a text corpus based on the vicuna-7b generated results on Stanford-alpaca, GSM8k and python-instruct-18k. However, this corpus is still not diverse enough, and also the text in it may deviate from the text generated by other LLMs, which limits the performance of SAM-Decoding. Therefore, in the future we need to collect more specialized and diverse corpus for different types of tasks.
On the other hand, when combining SAM-Decoding with other types of methods, we use a very heuristic approach, i.e., we choose different methods depending on the match length. This does not fully utilize the exact match lengths provided by the suffix automaton, so subsequently we will try to train classifier to select different decoding methods at each generate round.
Finally, the performance of retrieval-based methods is highly correlated with the usage scenarios, and the existing datasets do not well reflect the performance of retrieval-based methods in real usage, so in the future we also need to construct datasets that are more compatible with real scenarios to evaluate the performance of retrieval-based methods.
@@ A.1 Construction Process of Suffix Automaton
Algorithm 2 introduces the construction (Build-SAM) and expansion process (Expand) of Suffix Automaton, where the INIT_SAM function will create a suffix automaton that only contains the root node. For the root node, the link attribute value is $-1$, the next attribute value is empty, the length attribute value is 00, and the min_endpos attribute value is 00. Meanwhile, Algorithm 3 shows the construction process of the top-k successors for each node of static suffix automaton. Each node in the algorithm involves a new variable, “freq”, which represents the frequency of occurrence of the corresponding substring for each node, and can be initialized at the time of constructing the suffix automaton, i.e., “freq” is initialized to 1 for nodes generated by expansion, and “freq” is initialized to 0 for nodes generated based on cloning.
@@
Algorithm 3 Construction Process of Top-k Successors and Transition Probabilities
  function dfs
     Input: state $s$
     for $t_{n},s_{n}\in s.\text{next}$ do
        $\text{dfs}(s_{n})$
        $s.\text{freq}=s.\text{freq}+s_{n}.\text{freq}$
     end for
     $s.\text{topk\_succs}=\text{TopK}_{\text{freq}}(s.\text{next})$
[…]
        $s.\text{topk\_prob}.\text{append}(s_{n}.\text{freq}/s.\text{freq})$
     end for
  end function
  function Init_topk
     Input: suffix automaton $S$
     $\text{dfs}(S.\text{root})$
  end function
A.2 Drafting via Prim’s Algorithm
Algorithm 4 introduces a drafting process based on Prim’s algorithm to find a maximum spanning tree. For static suffix automata, we can offline maintain the frequency of occurrence of the corresponding substring for each node. Therefore, based on the recorded frequency for each node in the automaton, we can calculate the top-k successors and corresponding transition probabilities, where the transition probability is calculated by dividing the frequency of occurrence of the target state by the frequency of occurrence of the current state.
Algorithm 4 Drafting via Prim’s Algorithm
  function Prim
     Input: suffix automaton $S$, state $s$, start token $t$
     $q=\text{PriorityQueue}()$
     $q.\text{push}(\{1.0,s,t\})$
     $d=[]$
[…]
        $p,\,s,\,t=q.\text{top}()$
        $q.\text{pop}()$
        $d.\text{append}(t)$
[…]
           $p_{\rm new}=p*p_{n}$
           $s_{\rm new}=s_{n}$
           $t_{\rm new}=t_{n}$
           $q.\text{push}({p_{\rm new},s_{\rm new},t_{new}})$
        end for
     end while
     Output: draft tree $d$
  end function
A.3 Time Complexity of State Transfer
@@
Next, we analyze the upper bound of $\hat{c_{i}}$. Each state transition involves moving through the link edge zero or more times, followed by a move through the next edge. Transitioning through the link edge incurs a cost of 1 but decreases the potential by at least 1. Conversely, transitioning through the next edge incurs a cost of 1 and increases the potential by 1. Consequently, the amortized cost $\hat{c_{i}}$ is bounded above by 2, leading to:
@@
Model Method MT Trans Sum QA Math RAG #MAT Tokens/s Speedup
Llama3-8B PLD 1.30$\times$ 1.12$\times$ 1.41$\times$ 1.03$\times$ 1.30$\times$ 1.53$\times$ 1.39 44.26 1.28$\times$
SAM-Decoding 1.59$\times$ 1.35$\times$ 1.50$\times$ 1.35$\times$ 1.54$\times$ 1.75$\times$ 1.72 52.35 1.51$\times$
@@
SAM-Decoding[T] 2.09$\times$ 1.93$\times$ 2.04$\times$ 1.82$\times$ 2.32$\times$ 2.12$\times$ 2.63 71.73 2.05$\times$
EAGLE-2 2.08$\times$ 1.95$\times$ 1.85$\times$ 1.80$\times$ 2.31$\times$ 1.87$\times$ 3.90 68.69 1.98$\times$
SAM-Decoding[E2] 2.36$\times$ 1.96$\times$ 1.98$\times$ 1.79$\times$ 2.32$\times$ 2.11$\times$ 3.92 72.47 2.08$\times$
@@
Model Method HumanEval HAGRID
#MAT Tokens/s Speedup #MAT Tokens/s Speedup
Llama3-8B PLD 1.30 42.39 1.18$\times$ 1.50 45.15 1.56$\times$
SAM-Decoding 2.06 64.38 1.79$\times$ 1.88 58.40 2.02$\times$
Token Recycling 2.93 71.49 1.99$\times$ 2.84 62.77 2.17$\times$
SAM-Decoding[T] 2.77 78.04 2.16$\times$ 2.70 66.76 2.30$\times$
EAGLE-2 4.74 85.58 2.37$\times$ 3.97 63.30 2.18$\times$
SAM-Decoding[E2] 4.76 91.50 2.54$\times$ 3.93 67.94 2.35$\times$
Table 4: Speedup of SAM-Decoding compared to the baselines on HumanEval and HAGRID.
Model Method MT Trans Sum QA Math RAG #MAT Tokens/s Overall
@@
SAM-Decoding 2.08$\times$ 1.26$\times$ 2.23$\times$ 1.53$\times$ 2.09$\times$ 1.89$\times$ 2.19 39.24 1.84$\times$
@@
EAGLE-2 3.10$\times$ 2.15$\times$ 2.58$\times$ 2.38$\times$ 3.19$\times$ 2.33$\times$ 4.42 56.06 2.63$\times$
@@
Model Method HumanEval HAGRID
#MAT Tokens/s Speedup #MAT Tokens/s Speedup
Vicuna-13B PLD 1.54 32.06 1.44$\times$ 1.90 43.38 2.15$\times$
SAM-Decoding 2.42 48.92 2.20$\times$ 2.21 41.93 2.08$\times$
Token Recycling 2.79 46.03 2.07$\times$ 2.90 40.97 2.03$\times$
SAM-Decoding[T] 2.79 50.87 2.28$\times$ 2.99 48.33 2.40$\times$
EAGLE-2 5.15 77.85 3.49$\times$ 4.24 52.28 2.59$\times$
SAM-Decoding[E2] 5.12 78.96 3.54$\times$ 4.41 56.17 2.78$\times$
Table 6: Speedup of SAM-Decoding compared to the baselines on HumanEval and HAGRID.
Model Method MT Trans Sum QA Math RAG #MAT Tokens/s Overall
@@
SAM-Decoding 1.91$\times$ 1.25$\times$ 1.98$\times$ 1.48$\times$ 1.83$\times$ 1.66$\times$ 1.97 15.35 1.68$\times$
Token Recycling 2.10$\times$ 1.84$\times$ 2.19$\times$ 1.88$\times$ 2.42$\times$ 1.92$\times$ 2.70 18.80 2.06$\times$
@@
EAGLE-2 3.29$\times$ 2.31$\times$ 2.73$\times$ 2.51$\times$ 3.65$\times$ 2.46$\times$ 4.06 25.86 2.83$\times$
SAM-Decoding[E2] 3.40$\times$ 2.25$\times$ 2.93$\times$ 2.43$\times$ 3.45$\times$ 2.54$\times$ 4.08 25.91 2.84$\times$
@@
Model Method HumanEval HAGRID
#MAT Tokens/s Speedup #MAT Tokens/s Speedup
Vicuna-33B PLD 1.58 14.18 1.51×1.51$\times$1.51 × 1.55 15.74 1.80×1.80$\times$1.80 ×
SAM-Decoding 2.05 19.08 2.03×2.03$\times$2.03 × 1.90 16.15 1.85×1.85$\times$1.85 ×
Token Recycling 2.64 19.64 2.09×2.09$\times$2.09 × 2.71 18.29 2.09×2.09$\times$2.09 ×
SAM-Decoding[T] 2.73 22.44 2.39×2.39$\times$2.39 × 2.60 19.74 2.26×2.26$\times$2.26 ×
EAGLE-2 3.53 28.18 3.00×3.$0$ 3.84 24.28 2.78×2.78$\times$2.78 ×
SAM-Decoding[E2] 3.61 29.56 3.14×3.14$\times$3.14 × 3.82 25.08 2.87×2.87$\times$2.87 ×
Table 8: Speedup of SAM-Decoding compared to the baselines on HumanEval and HAGRID.
Model Method MT Trans Sum QA Math RAG #MAT Tokens/s Overall
Vicuna-7B Token Recycling 2.08$\times$ 1.76$\times$ 1.97$\times$ 1.85$\times$ 2.35$\times$ 1.76$\times$ 2.82 98.39 1.96$\times$
@@
EAGLE-2 2.66$\times$ 1.76$\times$ 2.18$\times$ 2.03$\times$ 2.63$\times$ 1.97$\times$ 4.34 110.56 2.21$\times$
@@
Table 9: Speedup of SAM-Decoding on A800 GPU compared to the baselines on Spec-Bench.
@@ Appendix B Additional Experiment Results
In this section, we present the results of the experiment on Llama3-8B-instruct, Vicuna-13B-v1.3 and Vicuna-33B-v1.3.
Tables 3 and 4 present the speedup ratios of SAM-Decoding compared to baseline methods across the Spec-Bench, HumanEval, and HAGRID datasets, utilizing the Llama3-8B-instruct model. It can be seen that the inference speed of SAM-Decoding outperforms the strongest retrieval-based baseline PLD on all tasks. Meanwhile, SAM-Decoding , when paired with Token Recycling (SAM-Decoding[T]), brings speedups on all tasks. Specifically, SAM-Decoding enhances the speedup ratio of Token Recycling from 1.92$\times$, 1.85$\times$, and 1.82$\times$ to 2.09$\times$, 2.04$\times$, and 2.12$\times$ for Multi-turn Conversation, Summarization, and Retrieval-Augmented Generation tasks, respectively. This improvement raises the overall speedup ratio of token recycling in the Spec-Bench dataset from 1.91$\times$ to 2.05$\times$. On the HumanEval and HAGRID datasets, SAM-Decoding increases the speedup ratio of Token Recycling from 1.99$\times$ and 2.17$\times$ to 2.16$\times$ and 2.30$\times$, respectively. Furthermore, SAM-Decoding also amplifies the performance gains of EAGLE-2 in Multi-turn Conversation, Summarization, Retrieval-augmented Generation, Code Generation and Context Q&A tasks. The speedup ratios were increased from 2.08$\times$, 1.85$\times$, 1.87$\times$, 2.37$\times$, and 2.18$\times$ to 2.36$\times$, 1.98$\times$, 2.11$\times$, 2.54$\times$ and 2.35$\times$ respectively.
Tables 5, 6, 7 and 8 present the speedup ratios of SAM-Decoding compared to baseline methods across the Spec-Bench, HumanEval, and HAGRID datasets, utilizing the Vicuna-13B-v1.3 and Vicuna-33B-v1.3. On both models, SAM-Decoding still has inference speed exceeding the retrieval-based baseline, while by combining Token Recycling and EAGLE-2 also further improves the inference speed of the model on the Multi-turn Conversation, Summarization, Retrieval-augmented Generation and Context Q&A tasks.